OpenStudy (vortish):

x=5-y complet the ordered pairs x =0,1,5

OpenStudy (anonymous):

so are you trying to find the y when x=0,1, and 5?

OpenStudy (vortish):

yes

OpenStudy (anonymous):

well you plug in each of those values for x and then solve for y

OpenStudy (anonymous):

so if x=5 then y =0 if x=1 then y=4 and if x=0 then y=5 right?

OpenStudy (vortish):

would those not be negetive

OpenStudy (anonymous):

no for example 1=5-y y has to be 4 cause if it were -4 then the equation would read 1=5-(-4) which is not true cause 9=5-(-4)